OpenStudy (aonz):

Help please :) Yr11 3 unit question on trig In triangle ABC, acosA = bcosB. Prove using the cosine rule, that the triangle is either isosceles or right-angled

OpenStudy (ajprincess):

Do u knw cos rule?

OpenStudy (aonz):

yep...

OpenStudy (ajprincess):

good:) so vat is cosA?

OpenStudy (aonz):

b^2 + c^2 - a^2 / 2bc

OpenStudy (ajprincess):

cosB=?

OpenStudy (aonz):

a^2 +c^2 - b^2 / 2ac

OpenStudy (ajprincess):

good:)

OpenStudy (ajprincess):

Nw if u r to prove that the triangle is isoscles u need to show that a=b or b=c or a=c. if u r to show it is a right angled triangle u should show pythagoras theorem is valid. \[a\cos A=b\cos B\] \[a* \frac{ b^2+c^2-a^2 }{ 2bc}=b*\frac{ a^2+c^2-b^2 }{ 2ac }\] vat do u get when u simplify this? @AonZ

OpenStudy (aonz):

2c^3 a^2 - 2a^4 c - 2c^3 b^2 + 2b^4 c = 0

OpenStudy (ajprincess):

u can factor out 2c.

OpenStudy (aonz):

2c ( b^4 - c^2 b^2 + a^2 c^2 - a^4) = 0 this right so far?

OpenStudy (ajprincess):

yup:) good:) nw divide by 2c both sides.

OpenStudy (aonz):

ok done that. btw im writing it down

OpenStudy (ajprincess):

ok:)

OpenStudy (ajprincess):

u will get this. b^4 - c^2 b^2 + a^2 c^2 - a^4=0 try to factorise this.

OpenStudy (aonz):

(b^2 - a^2)(b^2 +a^2 -c^2)

OpenStudy (aonz):

(b+a)(b-a)(b^2+a^2 - c^2)

OpenStudy (ajprincess):

(b+a)(b-a)(b^2+a^2 - c^2)=0 Divide by (b+a)(b-a) both sides. vat do u get?

OpenStudy (aonz):

b^2 +a^2 - c^2 = 0

OpenStudy (aonz):

you can tell its right angle by looking at this.. but how do you tell if its isosceles?

OpenStudy (ajprincess):

u r nt asked to prove both right? either one of that?

OpenStudy (aonz):

oh i guess :)

OpenStudy (aonz):

i still think it means to prove that the triangle must be right or isoscles but it cant be scalene

OpenStudy (aonz):

but can we show it isoscles anyways?

OpenStudy (ajprincess):

let me think. hav to go nw. will be back in 15 mnts

OpenStudy (anonymous):

(b+a)(b-a)(b^2+a^2 - c^2) =0 gives (b+a)=0 OR (b-a)= 0 OR (b^2+a^2 - c^2)=0 so, b-a = 0 --->a=b----->isosceles (b^2+a^2 - c^2)=0 --->right as you already know.

OpenStudy (aonz):

ohh xDDDD THANKS!! Shame i can only give medal to 1 person :(

OpenStudy (anonymous):

hey, i am not here for medals ;)

OpenStudy (ajprincess):

Welcome:) you can give it to @gohangoku58

OpenStudy (ajprincess):

cos he/she deserves it most:)

OpenStudy (anonymous):

:O You made him solve the whole thing and i deserve it!??

OpenStudy (aonz):

lol

OpenStudy (ajprincess):

yup:) it was u who pointed out the isoscles one:)